# FedSlate: A Federated Deep Reinforcement Learning Recommender System*

### v1

@@ VI Conclusion
To tackle the complexity of integrating user privacy data across diverse platforms into recommendation systems, we introduce a novel reinforcement learning algorithm, designated as FedSlate. This algorithm is designed to develop superior recommendation tactics collaboratively across multiple agents while safeguarding user privacy. Utilizing RecSim, we established a multi-platform recommendation simulation to assess how our algorithm benefits various participants within Federated Learning (FL). Our research indicates that FedSlate effectively resolves the challenges of cross-platform learning in recommendation systems without necessitating the exchange of private data between platforms. Looking ahead, we aim to refine and broaden the scope of FedSlate in the realm of cross-platform recommendation systems that prioritize privacy. To this end, we plan to augment FedSlate with cutting-edge privacy protection methods, such as secure aggregation, differential privacy, and FL with encrypted data, as detailed by Lyu et al. (2022)[47]. These enhancements will ensure the integrity of user privacy throughout the collaborative learning process, even when handling sensitive data. Moreover, we will undertake comprehensive experiments and evaluations of the FedSlate algorithm in authentic recommendation settings involving diverse platforms and user bases. This will yield critical insights into the algorithm’s efficiency, scalability, and practical applicability, alongside its influence on user satisfaction and engagement.
VII Acknowledgement
This work is supported by the National Natural Science Foundation of China, The Research Project of Shanghai Science and Technology Commission (Grant No.62102241, No.23ZR1425400).

### v2

@@ VI Conclusion
To tackle the complexity of integrating user privacy data across diverse platforms into recommendation systems, we introduce a novel reinforcement learning algorithm, designated as FedSlate. This algorithm is designed to develop superior recommendation tactics collaboratively across multiple agents while safeguarding user privacy. Utilizing RecSim, we established a multi-platform recommendation simulation to assess how our algorithm benefits various participants within FL. Our research indicates that FedSlate effectively resolves the challenges of cross-platform learning in recommendation systems without necessitating the exchange of private data between platforms.
VII Limitations
Although our FedSlate benefits all participating entities in federated recommendations, the introduction of FL incurs higher communication costs, potentially negating some of the performance enhancements in recommendation systems. Future research will investigate the thresholds for cost-benefit trade-offs. Additionally, the gains from FedSlate are unevenly distributed among the participants in the FL process. Entities lacking access to user feedback benefit disproportionately in terms of recommendation accuracy, receiving greater advantages. Subsequent efforts will focus on promoting fairness within the FL process by evaluating the contributions of each participant. Looking ahead, our objective is to refine and broaden FedSlate’s application in privacy-centric cross-platform recommendation systems. To this end, we aim to fortify FedSlate with advanced privacy-preserving technologies such as secure aggregation, differential privacy, and encrypted data in FL, thereby safeguarding user privacy throughout the collaborative learning process, even when handling sensitive data. Additionally, we will undertake extensive testing and evaluation of the FedSlate algorithm in real-world recommendation scenarios across diverse platforms and user demographics, yielding crucial insights into the algorithm’s efficiency, scalability, practical viability, and its effects on user satisfaction and engagement.
Biography Section
[Uncaptioned image] Yongxin Deng is a graduate student at the School of Electronic and Electrical Engineering, Shanghai University of Engineering Science, Shanghai, China. His research interests include artificial intelligence and healthcare informatics.
[Uncaptioned image] Xihe Qiu is an Associate Professor at the School of Electronic and Electrical Engineering, Shanghai University of Engineering Science, Shanghai, China. His research focuses on machine learning and clinical decision support systems. Email: qiuxihe@sues.edu.cn
[Uncaptioned image] Xiaoyu Tan received the B.E. degree in Mechanical Engineering from Northeastern University, China, in 2015, and the Ph.D. degree from the National University of Singapore, Singapore, in 2019. He worked as an Algorithm Engineer at Ant Group from 2019 to 2022. He is currently an AI Algorithm Engineer with INF Tech, Shanghai, China. His research interests include artificial intelligence and machine learning algorithms.
[Uncaptioned image] Yaochu Jin (Fellow, IEEE) received the BSc, MSc, and PhD degrees in automatic control from Zhejiang University, Hangzhou, China, in 1988, 1991, and 1996, respectively, and the Dr.-Ing. degree from Ruhr-University Bochum, Bochum, Germany, in 2001. He is presently Chair Professor of AI with the School of Engineering, Westlake University, Hangzhou, China. He is the Head of the Department of Artificial Engineering and leads the Trustworthy and General AI Laboratory. He was Alexander von Humboldt Professor for AI with the Faculty of Technology, Bielefeld University, Germany, and Surrey Distinguished Chair of Computational Intelligence, University of Surrey, U.K. He was a Finland Distinguished Professor and a Changjiang Distinguished Visiting Professor in China. His main research interests include data-driven evolutionary optimization, trustworthy machine learning, multiobjective evolutionary learning, and evolutionary developmental systems. Dr Jin is the recipient of the 2025 IEEE Frank Rosenblatt Award, the 2018, 2021 and 2024 IEEE Transactions on Evolutionary Computation Outstanding Paper Award, and the 2015, 2017, and 2020 IEEE Computational Intelligence Magazine Outstanding Paper Award. He is currently the President of the IEEE Computational Intelligence Society and Editor-in-Chief of Complex & Intelligent Systems. He is a member of Academia Europaea.